9, In the reaction 2O3 → 3O2(g) , the rate of appearance of Oxygen is 6.0 x 10-5 M/s at a certain moment. What will be the rate of disappearance of O3 at the same moment?


1
Expert's answer
2021-09-09T00:47:57-0400

We have to calculate rate of disappearance of O.

Rate = -1/2 ∆[O3]/∆t = 1/3 ∆[O2]/∆t 

We have ∆[O2]/∆t = 6.0*10-5 M/s 

So, 

-∆[O3]/∆t = 2/3 * ∆[O2 ]/∆t  

        = 2/3 * 6*10-5 ===> 4*10-5 M/s